OpenStudy (anonymous):

A triangular pyramid has a volume of 180 cm3 and a height of 12 cm. Find the length of a side of the triangular base if the triangle’s height from that side is 6 cm.

OpenStudy (anonymous):

(180x3)/12 = 45 (45x2)/6 = 15 the length of that side is 15cm